OpenStudy (anonymous):

Is the simplified form of 2square root of 3 - 2square root of 3 rational? Yes No

OpenStudy (jdoe0001):

\(\bf 2\sqrt{3}-2\sqrt{3}=?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so... what did you get?

OpenStudy (anonymous):

i dont know how to do it

OpenStudy (jdoe0001):

\(\bf 2\sqrt{3}\) looks a lot like \(\bf 2\sqrt{3}\) so... SAME - SAME = ?

OpenStudy (anonymous):

same ?

OpenStudy (jdoe0001):

hmm... 2 - 2 = 2?

OpenStudy (jdoe0001):

the same quantity subtracted from the same quantity

OpenStudy (anonymous):

no it = 0

OpenStudy (jdoe0001):

like say 4 -4, or 6-6 or 7-7 yeap, tis 0, and 0 is rational

OpenStudy (anonymous):

so its YES

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thank you so much

OpenStudy (nicia99):

Its no...